Of the whole numbers 10,11,,...,98,99, how many are greater than the sum if their digits?
A. 88
B. 89
C. 90
D. 99

Answer by jim_thompson5910(35256)

You can go through the entire list from 10 to 99 and you will not find any numbers where they are smaller than the sum of their digits.

So every number in this list matches the description. There are 99-10+1 = 89+1 = 90 numbers in that list

Final Answer: C) 90